# Efficient Response Generation Strategy Selection for Fine-Tuning Large Language Models Through Self-Aligned Perplexity

### v2

@@ 7 Conclusions
In this paper, we present a novel and scalable approach for selecting the optimal response generation strategy to train large language models. We introduce a new metric, self-aligned perplexity, which more effectively evaluates the alignment between the target model and the response options compared to traditional perplexity. Our experiments show that as the PGR increases, selecting the best response generation method becomes more crucial. We demonstrate that choosing the optimal generation strategy based on self-aligned perplexity can lead to substantial improvements in model performance, particularly on tasks with higher PGR. We hope our work will inspire researchers who use perplexity as a downstream metric and those seeking to generate the most effective responses from a set of inputs for training large language models.
@@ Appendix A Appendix
A.1 The Details of Strategy Selection Through Self-Aligned Perplexity
Perplexity
A straightforward way to gauge familiarity is via perplexity, which is typically defined as the exponentiation of the average negative log-likelihood per token. That is, for a candidate response consisting of tokens, we define:
In-Context Perplexity
To address this, we propose self-aligned perplexity, which incorporates ’s own initial prediction as an in-context example of style. Concretely:
1.
Generate responses, , where each response is generated from the corresponding input . may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation. To mitigate this, we employ a filtering mechanism to remove incorrect . According to our experimental results, since is primarily used to guide style, its correctness is less critical. Hence, the correctness filter is optional. For more details, please refer to Appendix A.4.
2.
For each candidate response , we assign in-context examples in a cyclic manner: the in-context example for the first input is based on the second input ’s generated response, the in-context example for is based on ’s response, and so on, until the in-context example for the last input is based on ’s response. This way, the model focuses on imitating the inference style rather than directly copying the answers.
3.
Construct a prompt, , containing both and a directive such as:
“Here is an example of how to reason and respond. Please strictly follow this style. If you drift, you must correct yourself mid-inference.”
The detailed prompt can be found in Appendix A.7.
4.
Compute perplexity for a candidate with the model forced to follow the self-demonstrated style in or .
Formally, define:
In this way, if is fluent but diverges from ’s reasoning style, the perplexity will be higher. This makes self-aligned perplexity a more reliable measure of style mismatch compared to plain token-level perplexity.
Selection Criterion.
For each strategy , we obtain from the small subset and average:
We then select:
@@ A.2 Can we get performance gain if we simply put all of the response variants together?
A.3 Spearman’s Rank Correlation Coefficient
Spearman’s rank correlation coefficient (Spearman’s ) is a non-parametric measure of rank correlation. Unlike Pearson’s correlation, which assesses the strength of a linear relationship between two continuous variables, Spearman’s focuses on the monotonic ordering of items. This makes it robust to outliers and suitable for evaluating ranking tasks where exact values matter less than the relative order in which items appear.
Formally, given two sets of ranks and for the same set of items, Spearman’s is defined as:
Here, and represent the rank of the -th item in each respective ranking. The numerator captures the sum of squared rank differences. If approaches 1, the two rankings are highly positively correlated; if nears 0, there is no monotonic relationship; and if approaches -1, the two rankings are inversely related.
In our experiments, we use the gold-standard orderings provided by the datasets as and the predicted rankings generated by our model and the train-then-test baseline as . By comparing these rankings via Spearman’s , we directly evaluate how faithfully each method captures the underlying ordering of items, providing a valuable complement to other accuracy-based metrics.
A.4 The Effectiveness of the Correctness Filter for Initial Predictions from the Target LLM
Before computing the semantic similarity between the initial prediction of the target LLM and the response from the teacher model, we apply a correctness filter(made by Qwen2.5. For more details, you can read the Appendix A.7) to remove incorrect predictions from the target LLM. In other words, we only compare the semantic similarity between correct initial predictions from the target LLM and responses from the teacher model. For details about how we make the
How effective is the correctness filter?
According to Table 6, the correctness filter helps to improve both accuracy and weighted spearman correlation.
| Weighted Method | Model | Accuracy | Weighted Spearman’s |
| Ours with filter | Mistral7B | 0.582 | 0.411 |
| Ours without filter | 0.574 | 0.364 | |
| Ours with filter | Llama3 | 0.650 | 0.344 |
| Ours without filter | 0.651 | 0.303 | |
| Ours with filter | Qwen2.5 | 0.723 | 0.494 |
| Ours without filter | 0.714 | 0.432 |
A.5 The Impact of Accuracy of the Synthetic Data on Training Outcomes
@@
A.6 Data Selection Rationale for the Benchmark
@@
A.7 Correctness Filter
Without supervised fine-tuning (SFT), may generate incorrect responses, making cosine similarity calculations between and unreliable. To alleviate this, we introduce a filtering mechanism to filter out the incorrect . We notice that for mathematical problems, the correct final answer typically appears as the last number in . Therefore, for Math-related tasks, we use regular expressions (regex) to extract the last number from the prediction and compare it directly with the ground truth. For other types of problem, we use the Qwen2.5-Instruct 7b model to extract the predicted label from the model output. We then compare this extracted label with the true gold label; if they match, we consider the prediction correct by default. The code we used to extract labels are detailed in Appendix A.8
@@ A.7 Correctness Filter
We construct redundant prompts to demonstrate that the perplexity of the redundant target responses is lower than that of GPT-4’s answers. Perplexity primarily reflects how fluent the language is and how well the language style aligns with the model, but it places less emphasis on semantic meaning.
Declaration of Independence
This is the part of Declaration of Independence that we use in the experiment in Table 4.
Self-Aligned In-Context Prompt for Perplexity Calculation This prompt shows how we add self - generated initial predictions from other questions as in - context examples for perplexity calculation.
A.7.1 Data Generation Strategies
@@
The prompt below is designed to guide GPT-4/Claude 3.5 in generating responses without relying on the ground truth solutions:
@@
The subsequent prompt guides GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Response (Rewrite GT) output.
@@
The following prompt directs GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Step-by-Step Response (No GT) responses.
@@
A.8 Prompt Used to Extract Labels from Predictions
The code below shows how we use Qwen-2.5-Instruct to extract the predicted labels from the predictions.
A.9 AI writing assistant
We use GPT-4o for writing assistant.
@@ A.9 AI writing assistant
| Gold Label | Mistral | ECQA | 0.383 | 0.181 | 0.722 | 0.251 | 0.084 | 0.562 |
| GPT-4o Answer Directly | 0.484 | 0.218 | 0.707 | 0.175 | 0.016 | 0.638 | ||
| Gold Label | Mistral | SQUAD | 0.082 | 0.0931 | 0.633 | 0.74 | 0.208 | 0.566 |
| GPT-4o Answer Directly | 0.512 | 0.234 | 0.594 | 0.748 | 0.268 | 0.628 | ||
| Gold Label | Mistral | DROP | 0.076 | 0.097 | 0.621 | 0.561 | 0.628 | 0.578 |
| GPT-4o Answer Directly | 0.542 | 0.241 | 0.602 | 0.546 | 0.736 | 0.638 | ||
| Gold Label | Mistral | WINOGRANDE | 0.381 | 0.172 | 0.625 | 0.166 | 0.042 | 0.742 |
| GPT-4o Answer Directly | 0.477 | 0.219 | 0.569 | 0.106 | 0.016 | 0.713 | ||
| Gold Label | LLAMA3 | ECQA | 0.798 | 0.416 | 0.734 | 0.193 | 0.1 | 0.637 |
| GPT-4o Answer Directly | 0.778 | 0.469 | 0.723 | 0.389 | 0.284 | 0.638 | ||
| Gold Label | LLAMA3 | SQUAD | 0.584 | 0.366 | 0.712 | 0.758 | 0.49 | 0.639 |
| GPT-4o Answer Directly | 0.791 | 0.457 | 0.726 | 0.759 | 0.368 | 0.651 | ||
| Gold Label | LLAMA3 | DROP | 0.144 | 0.169 | 0.674 | 0.574 | 0.738 | 0.582 |
| GPT-4o Answer Directly | 0.776 | 0.507 | 0.703 | 0.555 | 0.786 | 0.626 | ||
| Gold Label | LLAMA3 | WINOGRANDE | 0.776 | 0.445 | 0.717 | 0.226 | 0.162 | 0.766 |
| GPT-4o Answer Directly | 0.775 | 0.485 | 0.721 | 0.305 | 0.238 | 0.695 | ||
| Gold Label | Qwen | ECQA | 0.914 | 0.903 | 0.814 | 0.662 | 0.008 | 0.675 |
| GPT-4o Answer Directly | 0.903 | 0.888 | 0.793 | 0.668 | 0.016 | 0.716 | ||
| Gold Label | Qwen | SQUAD | 0.899 | 0.892 | 0.784 | 0.768 | 0.056 | 0.693 |
| GPT-4o Answer Directly | 0.896 | 0.911 | 0.789 | 0.756 | 0.074 | 0.712 | ||
| Gold Label | Qwen | DROP | 0.788 | 0.904 | 0.799 | 0.701 | 0.664 | 0.711 |
| GPT-4o Answer Directly | 0.911 | 0.903 | 0.792 | 0.741 | 0.806 | 0.701 | ||
| Gold Label | Qwen | WINOGRANDE | 0.893 | 0.904 | 0.78 | 0.651 | 0.004 | 0.725 |
| GPT-4o Answer Directly | 0.902 | 0.896 | 0.798 | 0.68 | 0.022 | 0.721 |

### v3

@@ 7 Conclusions
In this paper, we present a novel and scalable approach for selecting the optimal response-generation strategy to train large language models. We introduce a new metric, self-aligned perplexity, which more effectively evaluates the alignment between a target model and its response options compared to traditional perplexity. We demonstrate that choosing the optimal generation strategy based on self-aligned perplexity leads to substantial improvements in model performance, particularly on tasks with high performance variance. We hope our work will inspire researchers who use perplexity as a downstream metric or who wish to build the most effective instruction tuning datasets.
8 Limitations
While recent open-source “thinking” models(such as DeepSeek-R1-Distill-Qwen-7B) support long chain-of-thought reasoning, our evaluations focus on chat models. We used meta prompts to elicit reasoning steps but did not test on “thinking” models. We believe our style-alignment approach still applies, though further validation on smaller or differently pretrained models is needed. We leave broader scaling studies and extensions to other model families for future work.
@@ Appendix A Appendix
A.1 Ground Truth vs. Synthetic Data
As shown in Table 13 (Appendix), when ground truth is provided in natural language (e.g., GSM8K, MATH, ECQA, MBPP), training on ground truth is less effective than on synthetic data. This is because LLMs are more familiar with LLM-generated data, as demonstrated by Ren et al. (2024). However, when the ground truth is written as a gold label without a CoT inference process, training on the gold label can sometimes outperform training on CoT synthetic data within the same domain. However, in Table 11 (Appendix), training on gold labels harms cross-domain performance more than training on synthetic data. Besides, in real-life scenarios, training on natural language data is crucial, as users expect to see the rationale behind the final prediction made by LLMs.
@@ A.2 Can we get performance gain if we simply put all of the response variants together?
A.3 The Impact of Accuracy of the Synthetic Data on Training Outcomes
@@
A.4 Weighted Spearman’s Rank Correlation Coefficient
Spearman’s rank correlation () measures how well two orderings agree, ignoring absolute values. Because some of our {model, dataset} pairs exhibit far larger performance gaps among response-generation strategies than others, we assign higher importance to pairs whose choice of strategy matters more. We therefore adopt a weighted variant of Spearman’s correlation in which each item is given a non-negative weight .
Definition.
Let and be the ranks of the -th item under two orderings and let be its weight. Denote the weighted means
The weighted Spearman correlation is then the weighted Pearson correlation between the rank vectors:
Choice of weights.
For each {model, dataset} pair, we first train the target model on data produced by every candidate response-generation method and record the resulting accuracies. The weight is set to the standard deviation of these accuracies. Intuitively, tasks in which the strategies yield very different outcomes ( large) are more informative when judging a ranking metric, so they contribute more to .
Interpretation.
A value of indicates that the metric produces a ranking almost identical to the gold ranking, with higher-variance tasks influencing the score most strongly. Conversely, implies no weighted monotonic relationship, and signals an inverse agreement.
Throughout the main text and Appendix, all reported “Spearman” results actually correspond to this weighted formulation.
A.5 Data Selection Rationale for the Benchmark
@@
A.6 Correctness Filter
Without supervised fine-tuning (SFT), may generate incorrect responses, making cosine similarity calculations between and unreliable. To alleviate this, we introduce a filtering mechanism to filter out the incorrect . We notice that for mathematical problems, the correct final answer typically appears as the last number in . Therefore, for Math-related tasks, we use regular expressions (regex) to extract the last number from the prediction and compare it directly with the ground truth. For other types of problem, we use the Qwen2.5-Instruct 7b model to extract the predicted label from the model output. We then compare this extracted label with the true gold label; if they match, we consider the prediction correct by default.
| Methods | STD Range | #Data | Mistral | Llama | Qwen | Acc | Weighted |
| Ours - Ours w/o Filter | All | 51 | |||||
| Ours - Ours w/o Filter | 27 | ||||||
| Ours - Ours w/o Filter | 14 |
| Methods | STD Range | #Data | Mistral | Llama | Qwen | Acc | Weighted |
| Ours - Ours w/o Filter | All Data | 18.0 |
@@ A.6.1 Prompt for Self-Aligned Perplexity
We construct redundant prompts (shown in the Figure 7) to demonstrate that the perplexity of the redundant target responses is lower than that of GPT-4’s answers. Perplexity primarily reflects how fluent the language is and how well the language style aligns with the model, but it places less emphasis on semantic meaning.
Self-Aligned In-Context Prompt for Perplexity Calculation The prompt shown from the Figure 6 shows how we add self - generated initial predictions from other questions as in - context examples for perplexity calculation.
A.6.2 Data Generation Strategies
@@
The prompt from the Figure 2 below is designed to guide GPT-4/Claude 3.5 in generating responses without relying on the ground truth solutions:
@@
The subsequent prompt(shown in the Figure 5) guides GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Response (Rewrite GT) output.
@@
The following prompt from the Figure 4 directs GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Step-by-Step Response (No GT) responses.
@@
The following prompt from the Figure 3 directs GPT-4o to generate responses guided by GPT-4o generated Examples
@@
The following prompt from the Figure 3 directs GPT-4o to generate responses guided by Human written Examples
A.7 AI Assistant
We used GPT-4o as a writing assistant and programming aid for editing purposes.
A.8 Required Compute Resources
Each individual training run reported in this paper requires approximately 5–48 GPU hours when using a 40GB A100 GPU. We do not recommend you to reproduce every training run, as there are too many experiments. Instead, we strongly recommend directly using the reported training outcomes from each table as the final results. You can then compute your ranking metrics to evaluate how well your metric aligns with the training outcomes. Calculating metrics such as perplexity on a small subset of all of the dataset takes only about 2 hours on a single 40GB A100 GPU.
@@ A.8 Required Compute Resources
[…]
@@
| Data Generation Strategy | Model Type | plan bench generation | plan bench optimality | plan bench generalization | plan bench replaning | plan bench reuse | plan bench verification |
| gold label | mistral | 0.607 | 0.448 | 0.807 | 0.815 | 0.945 | 0.318 |
| groundtruth | |||||||
| gpt4 | 0.458 | 0.37 | 0.215 | 0.2 | 0.61 | 0.518 | |
| claude | 0.505 | 0.45 | 0.362 | 0.445 | 0.755 | 0.586 | |
| mini gpt4 | 0.445 | 0.38 | 0.282 | 0.235 | 0.51 | 0.5 | |
| step by step | 0.448 | 0.253 | 0.253 | 0.177 | 0.63 | 0.492 | |
| openai human written examples | 0.485 | 0.43 | 0.302 | 0.28 | 0.73 | 0.474 | |
| gpt4 style in context examples | 0.37 | 0.232 | 0.43 | 0.207 | 0.665 | 0.56 | |
| gold label | llama 3 instruct | 0.478 | 0.38 | 0.853 | 0.87 | 0.87 | 0.482 |
| groundtruth | |||||||
| gpt4 | 0.463 | 0.36 | 0.265 | 0.225 | 0.72 | 0.49 | |
| claude | 0.545 | 0.485 | 0.407 | 0.425 | 0.637 | 0.532 | |
| mini gpt4 | 0.415 | 0.375 | 0.25 | 0.215 | 0.5 | 0.674 | |
| step by step | 0.475 | 0.39 | 0.287 | 0.232 | 0.732 | 0.564 | |
| openai human written examples | 0.458 | 0.372 | 0.307 | 0.25 | 0.603 | 0.516 | |
| gpt4 style in context examples | 0.45 | 0.435 | 0.438 | 0.223 | 0.75 | 0.468 | |
| gold label | qwen | 0.32 | 0.285 | 0.38 | 0.732 | 0.845 | 0.516 |
| groundtruth | |||||||
| gpt4 | 0.237 | 0.253 | 0.145 | 0.2 | 0.515 | 0.538 | |
| claude | 0.325 | 0.265 | 0.277 | 0.46 | 0.393 | 0.534 | |
| mini gpt4 | 0.253 | 0.255 | 0.168 | 0.188 | 0.42 | 0.55 | |
| step by step | 0.265 | 0.182 | 0.135 | 0.17 | 0.577 | 0.536 | |
| openai human written examples | 0.24 | 0.19 | 0.222 | 0.155 | 0.47 | 0.52 | |
| gpt4 style in context examples | 0.302 | 0.268 | 0.338 | 0.2 | 0.597 | 0.534 |
A.9 License of the Dtasets
All dataset we use are publicly available dataset for research purpose. API-BANK (Lv 3 problems) (Li et al., 2023): CC-BY-SA GSM8K (Cobbe et al., 2021): MIT license PIQA (Bisk et al., 2020): unkown BoolQ (Clark et al., 2019):CC BY-SA 3.0 MBPP (Austin et al., 2021):CC BY 4.0 DROP (Dua et al., 2019): CC BY-SA 4.0 ARC-Challenge (Clark et al., 2018):CC BY-SA 4.0 PlanBench (Valmeekam et al., 2023): MIT license MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021): MIT license SQuAD (Rajpurkar et al., 2016):SA 4.0 license MMLU(Hendrycks et al., 2020): MIT license WinoGrande (Sakaguchi et al., 2021): Apache-2.0 license Hellaswag (Zellers et al., 2019): MIT license ECQA (Aggarwal et al., 2021): Apache-2.0 license MMLU_PRO (Wang et al., 2024): Apache-2.0 license AGIEval(Zhong et al., 2023): MIT license